# Joint Diffusion Models in Continual Learning

### v2

@@ 5.1 Experimental Setup
Metrics
To evaluate our approach, we use two metrics commonly used in continual learning: average accuracy and average forgetting, defined as in [6]. We average the evaluated metrics across 3 runs with different seeds and report the mean and standard deviation.
Implementation details
We use the same joint diffusion model architecture in all experiments on CIFAR10 and CIFAR100, and an upscaled version of it for ImageNet100 experiments to adjust to the bigger resolution of the dataset. For the UNet, we follow the implementation provided in [13]. For the classifier, we use a 2-layer feed-forward net with a hidden layer of 1024 and a LeakyRELU activation function. The detailed description of the training hyperparameters is provided in the Appendix, and in the code repository [^2] https://github.com/pskiers/Joint-Diffusion-in-Latent-Space.
5.2 Main results
To compare our method with other generative rehearsal approaches, we measure its performance on several CL benchmarks. In Tab. 1, we show that our method outperforms other evaluated generative replay techniques in terms of average accuracy after the last task. In particular, JDCL outperforms the recent state-of-the-art method–GUIDE on both average accuracy and forgetting. On CIFAR10 and ImageNet100, our method approaches the soft upper limit, calculated as a continual joint method trained with an infinitely long replay buffer. On those benchmarks, we surpass the previous state-of-the-art score by over 19 and 15 points, which is an increase of around 30% and 40% for CIFAR10 and ImageNet100 respectively.
Figure 3: Accuracy on each task after each phase of incremental training on CIFAR100 with 5 tasks
For CIFAR100 benchmarks, although JDCL outperforms other methods on the average accuracy metric, we observe higher forgetting when compared to the GFR. To better understand these results, we perform a detailed comparison of our method against the GFR and GUIDE methods. To that end, in Fig. 3 we present the accuracy of each approach, on each task after each training phase on CIFAR100 with 5 tasks. JDCL significantly outperforms GUIDE in terms of knowledge retention from preceding tasks. This is especially visible when we examine the change in performance of the methods on the older tasks. Whereas GUIDE’s knowledge obtained during the first task is quickly forgotten and is almost completely lost at the end of the training, JDCL is able to maintain a good accuracy on the first task for the whole training process. Notably, the performance degradation in our method seems to be related rather to the increasing complexity of the problem instead of forgetting, as we do not see fast degradation on the oldest tasks. On the other hand, when comparing to the GFR, we can observe that our method, though worse at preserving knowledge, is much better at assimilating new information. This phenomenon can be attributed to the significantly limited plasticity of the GFR that slows down the training of its feature extractor and restricts its update to the feature distillation loss. Consequently, the model has a substantially worse final performance.
Table 2: Average accuracy with a standard deviation of different methods tested with 5-task CIFAR-10 and 10-task CIFAR-100 semi-supervised settings. The number between brackets indicates the size of the memory buffer for the labeled data. Results of competing approaches from [25]
@@ 5.2 Main results
0.8% 5% 0.8% 5%
Fine-tuning 13.6$\pm$2.9 18.2$\pm$0.4 1.8$\pm$0.2 5.0$\pm$0.3
oEWC 13.7$\pm$1.2 17.6$\pm$1.2 1.4$\pm$0.1 4.7$\pm$0.1
ER (500) 36.3$\pm$1.1 51.9$\pm$4.5 8.2$\pm$0.1 13.7$\pm$0.6
iCaRL (500) 24.7$\pm$2.3 35.8$\pm$3.2 3.6$\pm$0.1 11.3$\pm$0.3
FOSTER (500) 43.3$\pm$0.7 51.9$\pm$1.3 4.7$\pm$0.6 14.1$\pm$0.6
X-DER (500) 33.4$\pm$1.2 48.2$\pm$1.7 8.9$\pm$0.3 18.3$\pm$0.5
PseudoER (500) 50.5$\pm$0.1 56.5$\pm$0.6 8.7$\pm$0.4 11.4$\pm$0.5
CCIC (500) 54.0$\pm$0.2 63.3$\pm$1.9 11.5$\pm$0.7 19.5$\pm$0.2
PAWS (500) 51.8$\pm$1.6 64.6$\pm$0.6 16.1$\pm$0.4 21.2$\pm$0.4
CSL (500) 64.5$\pm$0.7 69.6$\pm$0.5 23.6$\pm$0.3 26.2$\pm$0.5
NNCSL (500) 73.2$\pm$0.1 77.2$\pm$0.2 27.4$\pm$0.5 31.4$\pm$0.4
PseudoER (5120) 55.4$\pm$0.5 70.0$\pm$0.3 15.1$\pm$0.2 24.9$\pm$0.5
CCIC (5120) 55.2$\pm$1.4 74.3$\pm$1.7 12.0$\pm$0.3 29.5$\pm$0.4
ORDisCo (12500) 41.7$\pm$1.2 59.9 $\pm$1.4 - -
CSL (5120) 64.3$\pm$0.7 73.1$\pm$0.3 23.7$\pm$0.5 41.8$\pm$0.4
NNCSL (5120) 73.7$\pm$0.4 79.3$\pm$0.3 27.5$\pm$0.7 46.0$\pm$0.2
JDCL 78.93$\pm$0.72 79.96$\pm$0.86 22.19$\pm$0.3 26.39$\pm$1.7
Semi-supervised continual learning
Apart from the standard supervised setup, we evaluate our method on a set of semi-supervised continual scenarios and compare it to the existing approaches. As shown in Tab. 2 JDCL outperforms other methods on all CIFAR10 setups, while also achieving comparable accuracy on CIFAR100. What’s worth noting is that JDCL is able to reach state-of-the-art performance on CIFAR10 without a memory buffer, while also in a majority of tasks sampling fewer examples from the diffusion model than the buffer size in the related works. For CIFAR100, we believe that the reduced performance of our method is caused by an extremely weak signal from the classifier, which introduces an imbalance in our joint training. To further explain the issue, we visualize how JDCL build latent representations in the SSL setup in the appendix.
5.3 Additional Experiments
Ablation studies
In this work, we propose a new generative replay method for continual learning. Apart from the main idea of our approach which is joint diffusion and classification modeling, we combine it with two important techniques: knowledge distillation and two-stage training. In this section, we ablate the impact of each component. Notably, for experiments without joint modeling, we use a separate diffusion model with the same architecture as the joint diffusion model, and ResNet-18 as a classifier. We train and distill them in the same manner as the joint model. In table Tab. 3 we show the results for the model trained without individual components. As visible, the performance of our method is attributed to the synergy of the proposed joint modeling, knowledge distillation technique, and two-stage training.
Joint modeling Knowledge distillation 2-stage training Accuracy
83.7
68.4
48.2
36.7
32.7
JDCL as a continual representation learner
In this section, we show that our method is able to learn meaningful data representation in a continual setup. To that end, we follow the evaluation approach common in self-supervised continual learning techniques, as presented in [15]. In particular, we report the accuracy of a logistic regressor trained on the features provided by our model i.e. on the set of representations of the UNet model, pooled into a single vector. In Fig. 4 we compare our model with self-supervised continual methods. We can observe that the representations learned by our model are indeed very informative. Moreover, we find that JDCL outperforms other approaches on the CIFAR100 setup with 5 tasks. However, it is important to note that the other methods are self-supervised and thus, did not use labels during training.
Figure 4: Accuracy of a logistic regressor trained on representations learned with different self-supervised methods on CIFAR100. Results of competing approaches from [15]
Method CIFAR-10/5 CIFAR-100/5
FID Precision Recall FID Precision Recall
DGR 49.56 0.54 0.34 50.36 0.48 0.30
JDCL 32.12 0.60 0.43 39.93 0.71 0.27
Incremental generative setup comparison
Generative-replay-based continual learning methods rely heavily on the underlying generative model to preserve knowledge about the previous tasks. The quality of the samples generated by the method is therefore of utmost importance. In this section, we perform a comparison of JDCL and the diffusion model trained continually on CIFAR10 and CIFAR100, both split into 5 tasks. We present the achieved FID, Precision, and Recall after the last task in Table 4. JDCL outperforms the baseline method on all but one metric. We can observe significant improvement in the precision of generations, which results in generated rehearsal samples being clearly assigned to one of the previous classes bringing less noisy signals to the model. As argued in [18], this trait can be the reason for the improved performance of the continually trained classifier.
Runtime analysis of JDCL for CIFAR-10/5
Method Time [GPU-hours] Accuracy
DDGR $41.29$ 43.69
DGR diffusion 28.02 59.00
GUIDE 30.60 64.47
JDCL (fast) 27.21 78.10
JDCL 53.83 83.69
6 Conclusions

### v3

@@ Appendix C Stability of the JDCL’s representation space in an incremental setup
Appendix D Representation space in a semi-supervised setup
In this work, we introduce a new method of training joint diffusion model in a semi-supervised setup. In this section, we show that this approach allows our model to learn shared representations that are meaningful for both generative and discriminative objectives. To that end, we perform a visual comparison of the representation spaces in Fig. 8. We compare our method with a state-of-the-art semi-supervised method FixMatch [63], and a baseline classifier model that we train only on the labeled data. To extract the representations, we average pool the UNet representations into a single vector in our approach, while for the other two models, we take the output of the second to last layer. We observe that our approach is able to group the labeled and unlabeled data from the same class into similar regions of the embedding space shared between the classifier and diffusion model, similar to the state-of-the-art FixMatch approach.
Figure 9: Visualization of the Joint Diffusion with attention-based classifier. The classification method processes the representations from the shallowest to the deepest levels of UNet architecture using the attention mechanism to create a feature vector.
Table 7: Comparison of the classification accuracy and FID scores of the Joint Diffusion model trained with the attention-based and with the pool-based classifier.
@@ Appendix D Representation space in a semi-supervised setup
all labels 1000 labels 250 labels all labels 10000 labels 2500 labels
JD: Attention - acc 95.7% 93.9% 90.9% 78.3% 72.9% 63.4%
JD: Pooling - acc 96.4% 92.1% 86.6% 77.6% 69.9% 60.8%
JD: Attention - FID 15.9 28.2 28.5 23.1 - -
JD: Pooling - FID 7.9 27.0 27.4 17.4 - -
Appendix E Additional results - average forgetting
To further evaluate our approach, we use another metric commonly used in continual learning - average forgetting, defined as in [6]. In Tab. 6 we present the results of JDCL alongside competing generative replay techniques. Our method outperforms all other approaches on both CIFAR10 and ImageNet100. For CIFAR100 benchmarks, although JDCL outperforms other methods on the average accuracy metric, we observe higher forgetting when compared to the GFR. However, as discussed in Sec. 5.2 and shown in Fig. 3, this can be attributed to the limited plasticity of the GFR that significantly restricts its model’s update. This approach, while admittedly leading to lower average forgetting, results in worse plasticity, and consequently significantly worse final performance.
Appendix F Evaluation of generative qualities
To further evaluate the generative part of our method, we perform additional experiments. In Tab. 8, we compare the FID scores after every task of the CIFAR100/5 setup. Our method outperforms DGR (and GUIDE) after each task. To better understand the advantage of our approach, in Tab. 9 we calculate the equivalents of Forward and Backward Transfer (as in DER [79] for the FID metric. We observe that our method is significantly more stable and slightly less plastic, which is consistent with the results on the discriminative task. Nevertheless, JDCL’s overall FID performance is much higher, approaching the IID upper bound equal to 29.7.
Table 8: Comparison of the FID scores of JDCL and Deep Generative Replay with standard DDPM after each task of the CIFAR100/5 setup.
Method Task
1 2 3 4 5
DGR/GUIDE 19.08 29.52 39.11 40.75 50.36
JDCL 18.20 28.06 29.73 31.39 39.93
Table 9: Forward and backward FID transfer of JDCL and Deep Generative Replay on CIFAR100/5.
Method Forward Transfer ($\downarrow$) Backward Transfer ($\downarrow$)
DGR 19.80 10.22
JDCL 22.80 7.80
Appendix G Other ways to perform classification in the Joint Diffusion model
We extensively explored the design of the classifier during our preliminary studies. Here, we include a comparison of the selected pooling method with one of those alternative approaches - an attention-based classifier. Instead of pooling features into a single vector, this method applies a chain of attention blocks over the UNet’s h-representations. Each block takes the corresponding layer’s tensor and the previous block’s output to compute inter‑level attention, merging them into a single vector. Consequently, the classifier better exploits the representations and achieves higher accuracy on most tested setups, as shown in Tab. 7. However, training with the attention-based approach degrades the quality of the generations, making it less suitable for CL.
Appendix H Additional Backward and Forward Transfer metrics
In Tab. 10, we present the additional Forward and Backward Transfer metrics for JDCL, GUIDE, and GRF on CIFAR100/5 setup. The calculated metrics support our claims from Fig. 3. JDCL significantly outperforms GUIDE on Backward and GRF on Forward Transfer.
Table 10: Comparison of Forward and Backward Transfer of JDCL, GRF and GUIDE on CIFAR100/5 setup.
Method Forward Transfer ($\uparrow$) Backward Transfer ($\uparrow$)
GUIDE -15.72 -26.86
GRF -49.99 -10.00
JDCL -16.61 -14.83
Appendix I Details on semi-supervised scenario
All experiments follow offline CL protocols, while our semi-supervised scenario builds on the NNCSL [26], designed specifically for SSL in CL. In the paper authors adapt various offline approaches to the SSL framework, and we adhere to their evaluation protocol. Due to an additional diffusion model, we extend JDCL training to 60k batch updates for CIFAR10.
Appendix J Training complexity - discussion
While joint training increases the complexity of a single DM training step, it eliminates the need to train a separate classifier. Consequently, our joint model has fewer parameters (56.3 M) than the separate classifier and DM combined (64.7 M). Additionally, because of the synergy of the generative and discriminative parts, our method converges faster than the standard DM. We show that in Tab. 11, where we compare training loss on the first task of CIFAR10/5 of the 2 methods after the same runtime.
Table 11: Comparison of the training diffusion loss of the jointly and separately trained diffusion model on the first task of CIFAR10/5.
Runtime Joint diffusion Separate Diffusion
1h 0.01785 0.01982
2h 0.01741 0.01844
3h 0.01691 0.01723
Appendix K Sensitivity of performance with respect to loss terms scales
For all experiments, we use exactly the same classification loss scales equal to $0.001$ and non-knowledge-distillation loss scale $\beta$ equal to $0.01$. However, to present a full picture, we present the influence of different $\beta$ values on the final score on CIFAR10/5. Notably, the majority of results are within statistical errors:
$\beta$ 0 0.001 0.005 0.01 0.05 0.1 0.5 1
[…]